# Language Models Learn Rare Phenomena from Less Rare Phenomena: The Case of the Missing AANNs

### v1

@@ 2.4 Ablations
Common to subsequent experiments (Section 4 and Section 5) is the fact that we hold out certain parts of the BabyLM corpus—parts that conform to a certain linguistic or statistical hypothesis. This creates a gap between the experience of LMs trained on these ablated versions of the corpus, and that of the LM trained on the full BabyLM data. To circumvent this issue, we up-sample non-hypothesis-conforming utterances in BabyLM after performing our ablations, in a manner such that the LM still ends up encountering the exact same amount of tokens in such cases. As a result, all our LMs see the same quantity of tokens, but differ in content.
@@ 3 Experiment 1: LMs learn about aanns without having seen a single instance
3.1 LMs learn about aanns…
@@
To investigate the extent to which LMs trained on BabyLM learn the aann construction, we measure their accuracy on our tests described in Section 2.2. From Figure 2, we observe that the BabyLM-trained LMs obtain accuracies around 70%, which is substantially above chance. This suggests that LMs can reasonably acquire generalizations to aanns from exposure to positive evidence that makes up only 0.02% of their training experience.
@@
For comparison to larger, state-of-the-art LMs, we test Llama-2-7B (Touvron et al., 2023) and GPT-2 XL (Radford et al., 2019) on the aanns, and find them to achieve 83% and 78%, respectively. [^3] Qualitatively, performance is on these models is worse than data reported in Mahowald (2023), which used GPT-3. Because log-probabilities are not available from GPT-3, a direct comparison is not possible. Similarly, as a comparison to shallower LMs, we tested on 2-and 4-gram LMs trained on BabyLM and found both of them to achieve 0% accuracy, eliminating the possibility that the observed results are due to surface level statistics.
3.2 …without having seen a single instance…
Given that BabyLM-trained LMs learn the aann construction, how well would an LM generalize to aanns without having seen a single positive instance? To this end, we compare accuracies in the previous experiment to that obtained by LMs trained on BabyLM with our 2,301 detected aanns removed (i.e., no aann).
@@
From Figure 2, we find LMs trained with the no aann condition to achieve an average accuracy of 54%, which is a noticeable drop compared to the 70% obtained by the LMs trained on the complete BabyLM corpus, but importantly 47.75 points above chance performance (and, as we show below, well above comparable baselines with other constructions). This is a non-trivial result, since it suggests that LMs are able to learn the acceptability of a construction’s instances, without having seen a single positive occurrence, which indicates that there exist some systematic patterns in the corpus that are driving this generalization behavior.
@@
3.3 …more strongly than they learn counterfactual aann variants
To further contextualize the above results, we consider two counterfactual cases, where we replaced aanns in BabyLMs with instances involving the same lexical items, but in a word order that violates English grammar: (1) anan (e.g., a 90 whopping LMs); and (2) naan (e.g., 90 whopping a pages). This allows us to test if the results before are genuinely because LMs recognize the nuances of the aann construction, or whether they can learn alternative variants of the construction, which likely share fewer properties with other more frequent utterances in the data. If LMs are able to learn these constructions just as well as the the LMs in the previous experiments learned aanns, then the generalization claims from the previous experiments would be weakened.
To test for such possibilities, we create counterfactual versions of the Mahowald (2023) stimuli, where we apply analogous corruptions to the counterfactual variants of aann, such that they are violated in a similar manner as in the aann tests. Examples for the three types of instances in our tests can be found in Table 1. We then evaluate the previous two LMs (trained on BabyLM with and without seeing any aanns) with LMs trained on BabyLM with these counterfactual variants on judging the acceptability of aanns, anans, and naans. Figure 2 shows these results.
@@
From Figure 2, we make two high-level observations. First, and most importantly, LMs that see anans and naans do not learn those constructions as well as they learn aanns—especially the LM that saw no aanns (54% aann accuracy compared to 36% naan accuracy obtained by the naan-trained LM). Second, these LMs end up learning aanns more strongly relative to how well they learn constructions that they observe in lieu of the aann—e.g., naan trained LM achieves around 54% accuracy on aanns even though naans appeared frequently in the data set and no aanns did. This, combined with the results of the previous two sub-experiments suggests strongly that LMs pick up on cues from other—likely related—constructions encountered during training in order to assign non-trivial probability to unseen instances of aanns.
Figure 3: slors on aanns from Mahowald (2023) for our LMs (left) and a 4-gram baseline (right) trained on BabyLM and its ablated versions. Our LMs show a range of hypothesized effects, whereas the 4-gram baselines only show an effect of the adjective/numeral balanced condition. slor for unablated BabyLM-trained LM shown with dotted line.
@@
AANN a fine eighteen months 2,301
@@
6 months is a long time 62,597
@@
4 Experiment 2: Keys to Learning aanns
Our previous experiments reveal that, keeping everything else the same, LMs learn the aann construction far more accurately than they do its counterfactual variants. Furthermore, we also see strong aann acceptability judgments in LMs that have never encountered a single instance. This suggests that there are instances in the training data that contribute to the learning of the construction.
What might these be? Below we enumerate four hypotheses, each of which tackles subtly different aspects of the aann construction, then measure the effect of these phenomena, by separately holding them out during training and computing the average slor of the well-formed instances of the aann tests. The effect of a particular phenomenon on the acceptability of aanns can therefore be measured by comparing slors before and after holding out instances of that phenomenon. Methods for detecting the hypothesized phenomena can be found in Appendix B. As control, we additionally also hold out a random set of utterances, which do not conform to the target phenomena of interest. Note again that for all these cases, we ensure the LMs see the same number of tokens during training, by up-sampling other, non-aann containing sentences. Table 2 lists the hypotheses we consider, along with an example of their utterance and frequency of occurrence, in the BabyLM corpus.
The presence of “the ANN”
Phrases like “the beautiful five days” are common in corpora, which are not as unusual because “the” regularly takes plural nouns. We hypothesize that the acceptability of these structures affects the acceptability of aanns, since an LM might analogize from the general observation that ‘a’ or ‘an’ can substitute ‘the’ (e.g., a ball vs. the ball). Therefore, we consider all cases wheere a determiner precedes the contiguous sequence of article, numeral, plural noun.
A few/couple/dozen/etc. NNS
Another related phenomenon that is more common relative to the aann construction involves phrases such as “a few days” or “a couple bottles”. To an LM learner, they might provide evidence that measure noun phrases with plural nouns can be attached to an indefinite article (a/an), as is the case in aanns.
Measure NNS treated as singular
We consider yet another phenomenon involving phrases that treat measure nouns as singular, this time in terms of agreement, e.g., “Five miles is a long way to go”, and “1,000 pages is a lot for a dissertation.” These cases might provide further evidence to the model that measure noun phrases with plural nouns can be treated as a singular unit, thereby affecting the acceptability of the aann. These are less prevalent compared to the cases involving a few/couple/dozen NNS, but still far more frequent than the aann, therefore, we combine the two as a general case of treating measure NPs as singular.
@@ 4 Experiment 2: Keys to Learning aanns
A potential confound in the above ablations could be that the slor values of the aanns go down merely due to loss of content—this could be despite the fact that we add back additional tokens from BabyLM (such that all LMs see the exact same amount of tokens). To account for this, we additionally consider a control condition where we remove as many tokens as in the largest ablation (i.e., the A/An + ADJ/NUM case) such that none of the above phenomena are taken out. Insofar as the LMs rely on the aforementioned phenomena, results on LMs trained with this ablation should be closer to the original BabyLM-trained LM’s results.
@@ 4.1 Analysis and Results
In our experiments, we individually ablate out each of the aforementioned phenomena under two settings: (1) aanns are removed during training in addition to the phenomena; and when possible, (2) aanns are seen during training. (1) is a stricter setting, since here the LMs see neither the target phenomenon nor a single instance of the aann. Comparing average slors obtained in this condition to that obtained for the no aann can shed light on the the extent to which the target phenomenon is critical in allowing LMs to assign non-trivial probabilities on unseen aanns, zero-shot. On the other hand, (2) still allows for the LMs to perform lexical generalization (Kim et al., 2022) from seen aann instances—where they may arrive at strong acceptability measures on the test aanns by performing slot filling, without necessarily relying on the hypothesized phenomena (characterization of which is equally important).
@@
Figure 3 shows the average slors obtained across various ablations under the two settings. As a baseline, we compare our results to 4-gram LMs [^4] Trained using KenLM (Heafield, 2011) trained on corresponding ablations of the BabyLM corpus. We observe that holding out all our hypothesized phenomena has non-trivial effects on our LMs’ ratings of unseen well-formed aanns, and that these effects are different for when aanns are seen during training, or are held out. When aanns are held out along with the phenomena, we see substantial drops in the average slor values assigned by LMs on unseen aanns relative to that assigned by LMs in the no aann condition. Specifically, balancing the frequency of adjectives and numerals following an article has the greatest effect, followed by the two cases where measure nouns are treated as singular, followed by the removal of all cases that involve determiner + adjective + numeral + noun sequences. This suggests that in the absence of even a single aann during training, these phenomena are critical for LMs to assign non-trivial likelihoods to aanns. Importantly, simply ablating large amounts of data cannot explain these results, since LMs trained on our controlled condition show higher slor values than in all our hypothesis-informed ablations. These patterns are absent in 4-gram LMs, suggesting that they do not arise as a result of shallow, surface statistics, with the slight exception of differences observed in the case of the article+adjective/numeral ablation. Overall, this is a noteworthy finding, since it indicates that LMs can demonstrate a completely novel phenomenon (aann) by relying on other related—and more frequent—phenomena.
@@
When aanns are seen during training, however, we observe LMs’ results on unseen aanns to show more similar slor values with respect to the the LMs trained on the unablated BabyLM corpus, although they are still significantly reduced in some cases (e.g., singular measure nouns).
We conclude that when LMs see evidence of the aann construction, they do learn from it. But key related phenomena where measure nouns are treated as singular do seem to show some notable effects even when AANNs are present, suggesting that these enable additional learning even when aanns are present.
@@ 5 Experiment 3: The Role of Variability
Results from the previous experiment highlight the importance of seen aanns in order for LMs to generalize to unseen instances. What properties of these seen instances make LMs generalize? More broadly, there is a longstanding question as to how the nature of the instances of a construction provided during learning affect its (partial) productivity (Goldberg, 2005, 2019). In the context of aanns, we consider the role of variability on the open slots of the construction as a factor that might play a role in LMs’ productivity on unseen instances. The idea that slot-variability could affect learnability is motivated by theoretical claims in usage-based linguistics (Bybee, 1995), as well as existing research on novel constructions (Suttle and Goldberg, 2011), morphological productivity (Baayen, 2009; O’Donnell, 2015), and inductive inferences in cognitive psychology (Osherson et al., 1990; Xu and Tenenbaum, 2007). The idea is that encountering a slot with a wide variety of lexical items serves as a cue that the slot is flexible.
@@
We hypothesize that instances of aanns that provide natural evidence of greater open-slot variability—i.e. evidence that many different adjectives, numerals, and nouns can go in their respective positions in the aann construction—would lead LMs to assign greater likelihood to unseen aanns. On the other hand, LMs that only encounter a restricted set of instances might overfit, and be more conservative in extending the coverage of possible aanns to novel combinations of the slot-fillers. To test this, we divided our set of 2,301 aann-containing utterances in the BabyLM corpus into two equal subsets—one that contained aanns which were individually highly frequent but restricted in the types of adjective/numeral/nouns, and the other where the aanns were individually less frequent, but showed more variability in those slots. We then trained LMs on the BabyLM corpus containing utterances involving either of these two cases. Figure 4 shows the resulting average slors obtained from this experiment, along with those obtained by LMs trained on unablated BabyLM and the no aann conditions.
@@
From Figure 4, we see that LMs that only saw aanns that were highly variable in their open-slots demonstrated slors that were comparable and sometimes greater than those obtained by LMs that saw all aanns. By contrast, LMs that only saw aanns with low variability were as good as LMs that never saw any aanns. Therefore, LMs were sensitive to the nature of range of fillers that went into the construction’s open slots, showing relatively greater productivity when they observed evidence that the slots were highly variable. This is compatible with our hypothesis that slot-variability might affect the extent to which LMs “permit” productive uses of a construction.
@@
Figure 4: slors on aanns from Mahowald (2023) for LMs trained on BabyLM with low and high variability in the observed instances of aann. slor for unablated BabyLM-trained LM shown with dotted line.

### v2

@@ 2.6 Ablations
Common to subsequent experiments (§ 4 and § 5) is the fact that we hold out certain parts of the BabyLM corpus—parts that conform to a certain linguistic or statistical hypothesis. This creates a gap between the experience of LMs trained on these ablated versions of the corpus, and that of the LM trained on the full BabyLM data. To circumvent this issue, we up-sample non-hypothesis-conforming utterances in BabyLM after performing our ablations, in a manner such that the LM still encounters the exact same number of tokens.
@@ 3 Experiment 1: LMs learn about aanns without having seen a single instance
LMs learn about aanns…
@@
To investigate the extent to which LMs trained on BabyLM learn the aann construction, we measure their accuracy on our tests described in § 2.4. From Fig. 2, we observe that the BabyLM-trained LMs obtain accuracies around 70%, which is substantially above chance. This suggests that LMs can reasonably acquire knowledge of aanns, even though they make up only 0.02% of training utterances.
@@
For comparison to larger, state-of-the-art LMs, we test Llama-2-7B (Touvron et al., 2023) and GPT-2 XL (Radford et al., 2019) on the aanns. They got 83% and 78%, respectively. As a comparison to shallower LMs, we tested on 2-and 4-gram LMs trained on BabyLM and found both got 0% accuracy, strongly suggesting that the observed results are not due to n-gram statistics.
…without having seen a single instance…
@@
Given that BabyLM-trained LMs learn the aann construction, how well would an LM generalize to aanns without having seen a single positive instance? To this end, we compare accuracies in the previous experiment to that obtained by LMs trained on BabyLM with our 2,448 detected aanns removed (i.e., no aann). From Fig. 2, we find LMs trained with the no aann condition to achieve an average accuracy of 47%, which is a noticeable drop compared to the 70% obtained by the LMs trained on the complete BabyLM corpus, but importantly 40.75 points above chance performance (and, as we show below, well above comparable baselines with other constructions). This is a non-trivial result, since it suggests that LMs can learn the acceptability of a construction without having seen a single positive occurrence, which indicates that there exist systematic patterns in the corpus driving this generalization behavior.
@@
…more strongly than they learn counterfactual aann variants…
@@
To further contextualize the above results, we consider two counterfactual cases, where we replaced aanns in BabyLMs with instances involving the same lexical items, but in a word order that violates English grammar: (1) anan (e.g., a 90 whopping LMs); and (2) naan (e.g., 90 whopping a pages). This allows us to test if the results before are genuinely because LMs recognize the nuances of the aann construction. If LMs are able to learn these counterfactual constructions just as well as the LMs in the previous experiments learned aanns, then the generalization claims from the previous experiments would be weakened. To test for such possibilities, we create counterfactual versions of the Mahowald (2023) stimuli, where we apply analogous corruptions to the counterfactual variants of aann, such that they are violated in a similar manner as in the aann tests. Examples for the three types of instances in our tests can be found in Table 1. We then evaluate the previous two LMs (trained on BabyLM with and without seeing any aanns) with LMs trained on BabyLM with these counterfactual variants on judging the acceptability of aanns, anans, and naans. Fig. 2 shows these results, from which we make two high-level observations. First, and most importantly, LMs that see anans and naans do not learn those constructions as well as they learn aanns—especially the LM that saw no aanns (47% aann accuracy compared to 37% naan accuracy obtained by the naan-trained LM). Second, these LMs end up learning aanns better than they learn counterfactual constructions that they observed in lieu of the aann—e.g., naan trained LM achieves around 43% accuracy on aanns even though naans appeared frequently in the data and no aanns did. This, combined with the results of the previous two sub-experiments suggests strongly that LMs pick up on cues from other—likely related—constructions encountered during training in order to assign non-trivial probability to unseen instances of aanns.
…even with artificially polluted data…
[…]
Figure 3: z-scored aann acceptability ratings from humans and LMs trained on corpora with (1) aanns removed (i.e., no aann); and (2) left unablated for aanns with ‘Human’ nouns in Mahowald (2023)’s dataset. Even with ablated models, we observe the predicted dispreference for stubbornly distributive adjectives in the aann. Full results in Fig. 7.
…in a way that extends to lexical constraints.
While we focused on overall structural properties of aanns, there are also idiosyncrasies to the construction that arise from lexical semantic constraints. For instance, in many aann sentences, people prefer quantitative adjectives such as mere and hefty to qualitative ones such as beautiful (Mahowald, 2023; Solt, 2007) and find “stubbornly distributive” adjectives (“*a blue five pencils”) completely unacceptable (Schwarzschild, 2011). Insofar as our models learn aanns, we also should expect them to learn these lexical constraints. To test this, we compared LMs’ slors to human acceptability judgments on all 3.4k instances in Mahowald’s data across different adjective and noun classes. We found LMs trained on the original, unmodified BabyLM corpus to pattern similarly to humans in their preference for lexical constraints affecting aanns. Interestingly, these patterns were unchanged for LMs trained with the no aann condition, conforming to our predictions. For instance, as seen in Fig. 3, both our models share the human preference for quantitative and qualitative adjectives in the aann, compared to stubbornly distributive adjectives. More detailed results on lexical constraints can be found in App. E and we hypothesize that our broader set of results extends to include learning of lexical constraints on the construction.
4 Experiment 2: Keys to Learning aanns
Experiment 1 reveals that, keeping everything else the same, LMs learn the aann construction more accurately than they do its counterfactual variants. Furthermore, we also see strong aann acceptability judgments in LMs that have (almost) never encountered a single instance. This suggests that there could be instances in the training data that contribute to the learning of the construction.
What might these be? Below we enumerate four hypotheses, each of which tackles subtly different aspects of the aann construction, and then measure the effect of these phenomena by separately holding them out during training and computing the average slor of the well-formed instances of the aann tests. The effect of a particular phenomenon on the acceptability of aanns can therefore be measured by comparing slors before and after holding out instances of that phenomenon. Methods for detecting the hypothesized phenomena can be found in App. C. As control, we additionally also hold out a random set of utterances, which do not conform to the target phenomena of interest. Table 2 lists the hypotheses we consider, along with an example of their utterance and frequency of occurrence, in the BabyLM corpus.
The presence of “the ANN”
Phrases like “the beautiful five days” are common in corpora, which are not as unusual because “the” regularly takes plural nouns. We hypothesize that the acceptability of these structures affects the acceptability of aanns, since an LM might analogize from the general observation that ‘a’ or ‘an’ can substitute ‘the’ (e.g., a ball vs. the ball). Therefore, we consider all cases where a determiner precedes the contiguous sequence of article, numeral, plural noun.
A few/couple/dozen/etc. NNS
Another related phenomenon that is more common relative to the aann construction involves phrases such as “a few days” or “a couple bottles”. To an LM learner, they might provide evidence that measure noun phrases with plural nouns can be attached to an indefinite article (a/an; Solt, 2007), as is the case in aanns.
Measure NNS treated as singular
We consider yet another phenomenon involving phrases that treat measure nouns as singular, this time in terms of agreement, e.g., “Five miles is a long way to go”, and “1,000 pages is a lot for a dissertation.” These cases might provide further evidence to the model that measure noun phrases with plural nouns can be treated as a singular unit (Solt, 2007), thereby affecting the acceptability of the aann. These are less prevalent compared to the cases involving a few/couple/dozen NNS, but still far more frequent than the aann, therefore, we combine the two as a general case of treating measure NPs as singular.
@@ 4 Experiment 2: Keys to Learning aanns
AANN a fine eighteen months 2,448
@@
6 months is a long time 62,744
@@
Figure 4: slors on aanns from Mahowald (2023) for our LMs (left) and a 4-gram baseline (right) trained on BabyLM and ablated versions. Our LMs show a range of hypothesized effects, suggesting they contribute to aann learning. In contrast, the 4-gram LMs show mostly null results (except for the adjective/numeral balanced condition, which is highly sensitive to n-gram frequencies). The dotted line is slor for an unablated BabyLM-trained LM.
@@
A potential confound in the above could be that the slor values of the aanns go down merely due to loss of content, even though we add back additional tokens from BabyLM (such that all LMs see the exact same amount of tokens). To account for this, we additionally consider a control where we remove as many tokens as in the largest ablation (i.e., the A/An + ADJ/NUM case) such that none of the above phenomena are taken out.
@@ 4.1 Analysis and Results
In our experiments, we individually ablate out each of the aforementioned phenomena under two settings: (1) aanns are removed during training in addition to the target phenomena; and when possible, (2) aanns are seen during training. (1) is a stricter setting, since here the LMs see neither the target phenomenon nor a single instance of the aann. Comparing average slors obtained in this condition to that obtained for the no aann can shed light on the extent to which the target phenomenon is critical in allowing LMs to assign non-trivial probabilities on unseen aanns, zero-shot. On the other hand, (2) still allows for LMs to perform lexical generalization (Kim et al., 2022), where they may exhibit strong probabilities on the test aanns by performing slot filling, without necessarily relying on the hypothesized phenomena.
@@
Fig. 4 shows the average slors obtained across various ablations under the two settings. As a baseline, we compare our results to 4-gram LMs, trained using KenLM (Heafield, 2011), on corresponding ablations of the BabyLM corpus. We observe that holding out most of our hypothesized phenomena has non-trivial effects on our LMs’ ratings of unseen aanns, and that these effects are different for when aanns are seen during training, or are held out. When aanns are held out along with the phenomena, we see substantial drops in the average slor values assigned by LMs on unseen aanns relative to that assigned by LMs in the no aann condition. Specifically, balancing the frequency of adjectives and numerals following an article, along with the two cases where measure nouns are treated as singular, have the greatest effect. This suggests that, in the absence of even a single aann during training, these phenomena are critical for LMs to assign probability to aanns. Interestingly, holding out cases that involve any determiner + adjective + numeral + noun sequence has almost no impact relative to LMs trained on a corpus without only the aanns removed. Simply ablating large amounts of data cannot explain these results, since LMs trained on our controlled condition show higher slor values than in our hypothesis-informed ablations. These patterns are absent in 4-gram LMs, suggesting that they do not arise as a result of shallow statistics—with the exception of differences observed for the article+adjective/numeral ablation. Overall, this finding indicates that LMs can demonstrate a novel phenomenon (aann) by relying on other related—and more frequent—phenomena.
@@
When aanns are seen during training, however, we observe LMs’ results on unseen aanns to show more similar slor values with respect to the LMs trained on the unablated BabyLM corpus, although they are still significantly reduced in some cases (e.g., singular measure nouns). We conclude that direct evidence of observing instances of aann construction substantially enables generalization to unseen instances. At the same time, the presence of some key related phenomena in addition to direct evidence has an additive effect on this generalization behavior.
@@ 5 Experiment 3: The Role of Variability
Results from Experiment 2 highlight the importance of seen aanns in order for LMs to generalize to unseen instances. What properties of these seen instances facilitate LMs generalization behavior? This broadly relates to a longstanding question as to how the nature of the instances of a construction provided during learning affect its (partial) productivity (Goldberg, 2005, 2019). In the context of aanns, we consider the role of variability on the open slots of the construction as a factor that might play a role in LMs’ productivity on unseen instances. Encountering a slot with a wide variety of lexical items could serve as a cue that the slot is flexible. The idea that instance-variability could affect learnability is motivated by theoretical claims in usage-based linguistics (Bybee, 1995), as well as existing research on novel constructions (Suttle and Goldberg, 2011), morphological productivity (Baayen, 2009; O’Donnell, 2015), and inductive inferences in cognitive psychology (Osherson et al., 1990; Xu and Tenenbaum, 2007).
@@
We hypothesize that instances of aanns that provide natural evidence of greater open-slot variability—i.e. evidence that many different adjectives, numerals, and nouns can go in their respective positions in the aann construction—would lead LMs to assign greater likelihood to unseen aanns. On the other hand, LMs that encounter only a restricted set of instances might be more conservative in extending the coverage of possible aanns to novel combinations of the slot-fillers. To test this, we divided our set of 2,448 aann-containing utterances in the BabyLM corpus into two roughly equal subsets—one that contained aanns which were restricted in which tokens occur in a particular slot (low variability), and the other where the aanns showed more variability in those slots. We obtain these subsets by performing a median split based on the number of unique occurrences in a target slot(s), which resulted in a set of 1224 low and high variability instances. We repeated this for all three open slots (adjective/numeral/noun) jointly as well as those slots individually—i.e., 4 different types of target slots and 2 conditions each (low vs. high variability). Details about the slot fillers and examples from each set are provided in App. F. We then trained LMs on the BabyLM corpus containing utterances involving either of these two cases. Fig. 5 shows the resulting average slors obtained from this experiment, along with those obtained by LMs trained on unablated BabyLM and the no aann conditions.
@@
We see that the slor patterns of LMs trained on corpora that differed in aann slot-variability lie between the slor values elicited by LMs that never saw aanns and ones that saw every single aann in the original corpus. Among these, LMs that saw aanns that were highly variable in their open-slots elicited slors that were greater than those elicited by LMs that saw aanns with low open-slot variability. This was true for all cases except when “Numeral” was the target slot, where both variability conditions resulted in roughly similar slors. (We hypothesize that numerals may pattern differently since they may be inherently more fungible than other word classes.) Overall, these results suggest that LMs are sensitive to the nature of range of fillers that go into the construction’s open slots, showing relatively greater productivity when they observe evidence that the slots were highly variable. This is compatible with our hypothesis that slot-variability might affect the extent to which LMs “permit” productive uses of a construction.
@@
Figure 5: slors on aanns from Mahowald (2023) for LMs trained on BabyLM with low and high variability in the open slots of the observed aann instances. When models are presented with higher variability for a given slot, the construction is typically learned better.